OpenStudy (anonymous):

What is the radius of of a circle with a circumference of 18π ft? A. 6 ft B. 9 ft C. 18 ft D. 9π ft Is It B Or D?

OpenStudy (anonymous):

Is It B Or D?

OpenStudy (wolf1728):

Circumference = 2 * PI * radius, So, 18 PI = 2 * PI * radius radius = 18 PI / (2* PI) I think you can solve it from there.

OpenStudy (anonymous):

So Its Not B Or D >.<

OpenStudy (wolf1728):

Calculate 18 * PI / 2 * PI

OpenStudy (mustbethestars):

Or 18 * 3.14 / 2 * 3.14

OpenStudy (anonymous):

88.7364?

OpenStudy (wolf1728):

Okay solve (18/2) * (PI/PI)

OpenStudy (anonymous):

9

OpenStudy (wolf1728):

Very good !!

OpenStudy (anonymous):

Thanks So Is It 9ft Or 9rr Ft?

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

Is It B Or D? @wolf1728

OpenStudy (anonymous):

@ChantySquirrel1129**

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@jabez177

TheSmartOne (thesmartone):

You got 9 as your final answer you didn't get 9pi as your final answer.

TheSmartOne (thesmartone):

π is not rr. It's called pi, and is approximately 3.14

jabez177 (jabez177):

It is a number that goes on forever and never repeats itself.

jabez177 (jabez177):

Yes.

OpenStudy (anonymous):

Thanks :)